Find the slope of the line that passes through the pair of points (4.5, 13.5) and (–3.5, –9).
a.
0.36
c.
–1.26
b.
2.81
d.
–0.4

Answer by jim_thompson5910(35256)

Change in y: -9 - 13.5 = -22.5

Change in x: -3.5 - 4.5 = -8

-------------------------------------------------------


Slope = (Rise)/(Run)

Slope = (Change in y)/(Change in x)

Slope = -22.5/(-8)

Slope = 2.8125